# Copresheaf Topological Neural Networks: A Generalized Deep Learning Framework

### v2

@@ Abstract
We introduce copresheaf topological neural networks (CTNNs), a powerful and unifying framework that encapsulates a wide spectrum of deep learning architectures, designed to operate on structured data: including images, point clouds, graphs, meshes, and topological manifolds. While deep learning has profoundly impacted domains ranging from digital assistants to autonomous systems, the principled design of neural architectures tailored to specific tasks and data types remains one of the field’s most persistent open challenges. CTNNs address this gap by grounding model design in the language of copresheaves, a concept from algebraic topology that generalizes and subsumes most practical deep learning models in use today. This abstract yet constructive formulation yields a rich design space from which theoretically sound and practically effective solutions can be derived to tackle core challenges in representation learning: long-range dependencies, oversmoothing, heterophily, and non-Euclidean domains. Our empirical results on structured data benchmarks demonstrate that CTNNs consistently outperform conventional baselines, particularly in tasks requiring hierarchical or localized sensitivity. These results underscore CTNNs as a principled, multi-scale foundation for the next generation of deep learning architectures.
@@ 1 Introduction
Deep learning has excelled by exploiting structural biases—convolutions for images (Krizhevsky et al., 2012), transformers for sequences (Vaswani et al., 2017), and message-passing for graphs (Gilmer et al., 2017). Yet, the design of architectures that generalize across domains with complex, irregular, or multiscale structure remains a notorious challenge (Bronstein et al., 2017; Hajij et al., 2023b). Real-world data—spanning physical systems, biomedical signals, and scientific simulations—rarely adhere to the regularity assumptions embedded in conventional architectures. These data are inherently heterogeneous, directional, and hierarchical, often involving relations beyond pairwise connections or symmetric neighborhoods.
@@
Convolutional networks, designed for uniform grids, gloss over local irregularities; graph neural networks (GNNs) often rely on homophily and oversmooth features as depth increases; and transformers, while excellent at capturing long-range dependencies, assume homogeneous embedding spaces, incur quadratic complexity, and lack built-in notions of anisotropy or variable local structures. These shortcomings underscore the need for a framework that can natively encode diverse local behaviors, respect directional couplings, and propagate information across scales without sacrificing critical distinctions or imposing artificial uniformity.
@@
To address this foundational gap, we propose copresheaf topological neural networks (CTNNs), a unifying framework for deep learning based on copresheaves—a categorical structure that equips each local region of a domain with its own feature space, along with learnable maps specifying how information flows across regions. Unlike traditional models that assume a global latent space and isotropic propagation, our framework respects local variability in representation and directional flow of information, enabling architectures that are inherently multiscale, anisotropic, and expressive.
@@
Figure 1: A copresheaf topological neural network (CTNN) operates on combinatorial complexes (CCs), which generalize Euclidean grids, graphs, meshes, and hypergraphs. A CTNN is characterized by a set of locally indexed copresheaf maps $\rho_{x_{i}\to x_{j}}$, defined between cells $x_{i}$ and $x_{j}$ in the CC, and directed from $x_{i}$ to $x_{j}$. The figure illustrates how a CTNN updates a local representation $\mathbf{h}_{x}$ of a cell $x$ using neighborhood representations $\mathbf{h}_{y}$ and $\mathbf{h}_{z}$, which are sent to $x$ via the learnable local copresheaf maps $\rho_{x\to x}$, $\rho_{y\to x}$ and $\rho_{z\to x}$.
@@
By constructing our CTNNs upon combinatorial complexes (Hajij et al., 2023b, a), which generalize graphs, simplicial complexes, and cell complexes, we enable a principled message-passing mechanism over general topological domains grounded in the theory of copresheaves. This unified perspective subsumes and extends many deep learning paradigms—including GNNs, attention mechanisms, sheaf neural networks (Hansen and Ghrist, 2019d; Bodnar et al., 2022a), and topological neural networks (Papillon et al., 2023a; Hajij et al., 2023b; Bodnar et al., 2021; Ebli et al., 2020; Giusti et al., 2022b) within a single formalism. Our approach further departs from the traditional assumption of a single shared latent space by modeling task-specific, directional latent spaces that bridge diverse deep learning frameworks. Furthermore, CTNNs flexibly handle both Euclidean and non-Euclidean data, supporting expressive architectures such as Copresheaf GNNs, Transformers, and Convolutional Networks, which learn structure-aware, directional transport maps. Overall, CTNNs offer a promising framework for developing next-generation, topologically grounded, structure-aware machine learning models. See Figure 1 for an illustration.
Related work
@@
Our work here is closely related to sheaf neural networks (SNNs) extend traditional graph neural networks by leveraging the mathematical framework of cellular sheaves to capture higher-order and heterogeneous relationships. Early work by Hansen and Ghrist (2019a, b) introduced methods to learn sheaf Laplacians from smooth signals and developed a spectral theory that connects sheaf topology with graph structure. Building on these ideas, Hansen and Gebhart (2020a) proposed the first SNN architecture, demonstrating that incorporating edge-specific linear maps can improve performance on tasks involving asymmetric and heterogeneous relations.
@@
Recent advances have focused on mitigating common issues in graph neural networks such as oversmoothing and heterophily. For instance, Bodnar et al. (2022c) introduced neural sheaf diffusion processes that address these challenges by embedding topological constraints into the learning process. Similarly, Barbero et al. (2022a, b) developed connection Laplacian methods and attention-based mechanisms that further enhance the expressiveness and efficiency of SNNs. The versatility of the sheaf framework has also been demonstrated through its extension to hypergraphs and heterogeneous graphs (Duta et al., 2023; Braithwaite et al., 2024), which enables modeling of complex, higher-order interactions. Moreover, novel approaches incorporating joint diffusion processes (Caralt et al., 2024) and Bayesian formulations (Gillespie et al., 2024) have improved the robustness and uncertainty quantification of SNNs. Finally, the application of sheaf neural networks in recommender systems (Purificato et al., 2024) exemplifies their practical utility in real-world domains. Together, these contributions underscore the potential of SNNs to enrich graph-based learning by integrating topological and geometric information directly into neural architectures. Our proposed CTNNs subsume these architectures while avoiding restrictive co-boundary maps or rank-specific Laplacian operators. The appendix provides a more thorough literature review of related work.
@@ 2 Preliminaries
This section presents preliminary concepts needed for developing our theoretical framework. It recalls combinatorial complexes and neighborhood structures, reviews sheaves and copresheaves on directed graphs, and contrasts cellular sheaves with copresheaves in graph-based modeling.
@@ 2.1 Combinatorial Complexes and Neighborhood Structures
To ensure utmost generality, we base our framework on the combinatorial complexes (CCs) (Hajij et al., 2023b, a), which unify set-type and hierarchical relations over which data are defined. CC-neighborhood functions then formalize local interactions forming an ideal ground to define sheaves and higher-order message passing schemes for our copresheaf topological neural networks (CTNNs).
@@
A combinatorial complex (CC) is a triple $(\mathcal{S},\mathcal{X},\operatorname{rk})$, where $\mathcal{S}$ is a finite non-empty set of vertices, $\mathcal{X}\subset\mathcal{P}(\mathcal{S})\setminus\{\emptyset\}$, with $\mathcal{P}(\mathcal{S})$ denoting the power set of $\mathcal{S}$, $\operatorname{rk}:\mathcal{X}\to\mathbb{Z}_{\geq 0}$ is a rank function such that $\{s\}\in\mathcal{X}$, $\operatorname{rk}(\{s\})=0$ for all $s\in\mathcal{S}$, and $x\subseteq y\implies\operatorname{rk}(x)\leq\operatorname{rk}(y)$ for all $x,y\in\mathcal{X}$.
@@
When context permits, we write a CC $(\mathcal{S},\mathcal{X},\operatorname{rk})$ simply as $\mathcal{X}$; each $x\in\mathcal{X}$ has rank $\operatorname{rk}(x)$, $\dim\mathcal{X}=\max_{x\in\mathcal{X}}\operatorname{rk}(x)$, its $k$-cells $x^{k}$ are defined to be the cells $x$ with $\operatorname{rk}(x)=k$, furthermore we set $\mathcal{X}^{k}=\{x\in\mathcal{X}:\operatorname{rk}(x)=k\}=\operatorname{rk}^{-1}(\{k\})$. See Fig 2 for an example.
@@
Given a CC $\mathcal{X}$, a neighborhood function on a CC $(\mathcal{S},\mathcal{X},\operatorname{rk})$ is a map $\mathcal{N}:\mathcal{X}\to\mathcal{P}(\mathcal{X})$, which assigns to each cell $x$ in $\mathcal{X}$ a collection of neighbor cells $\mathcal{N}(x)\subset\mathcal{X}$, referred to as the neighborhood of $x$. In our context, two neighborhood functions are common, namely the adjacency neighborhood and the incidence neighborhood: (i) the adjacency neighborhood $\mathcal{N}_{\mathrm{adj}}(x)=\{\,y\in\mathcal{X}\mid\exists\,z\in\mathcal{X}:x\subset z,\;y\subset z\}$ and (ii) the incidence neighborhood $\mathcal{N}_{\mathrm{inc}}(x)=\{\,y\in\mathcal{X}\mid x\subset y\}$, potentially defined with respect to particular ranks.
@@
Let $\mathcal{N}$ be a neighborhood function on a CC $\mathcal{X}$. Let $\mathcal{Y}=\{y_{1},\dots,y_{n}\}\subset\mathcal{\mathcal{X}}$, $\mathcal{Z}=\{z_{1},\dots,z_{m}\}\subset\mathcal{\mathcal{X}}$ be two collections of cells s.t. $\mathcal{N}(y_{j})\subseteq\mathcal{Z}$ for all $1\leq j\leq n$. An element of the neighborhood matrix $\mathbf{G}\in\{0,1\}^{m\times n}$ is defined as $[\mathbf{G}]_{ij}=\mathbbm{1}_{\mathcal{N}(y_{j})}\left(z_{i}\right)$.
@@ 2.2 Sheaves and Copresheaves on Directed Graphs
The copresheaf formalism assigns each vertex its unique, natural feature space $\mathcal{F}(x)$, respecting the potentially heterogeneous nature of the data, and each directed edge a transformation $\rho_{x\!\to\!y}$ that tells how data move between those spaces, $\mathcal{F}(x)\to\mathcal{F}(y)$. This separation of where data reside from how they travel provides a principled backbone for learning beyond the single-latent-space assumption of standard deep learning architectures. Concretely, copresheaves are defined as follows [^1] While we avoid overly complicated jargon, the appendix links our constructs to those of category theory for a more rigorous exposition.:
Definition 4 (Copresheaf).
@@
Think of a copresheaf as a system for sending messages across a network, where each node has its own language (stalk), and edges translate messages (linear maps) to match the recipient’s language. More concretely, on a directed graph $G=(V,E)$, each vertex $x\in V$ carries a task-specific latent space $\mathcal{F}(x)$, and every edge $x\!\to\!y\in E$ applies a learnable, edge-indexed linear map $\rho_{x\to y}:\mathcal{F}(x)\to\mathcal{F}(y)$ that re-embeds $x$’s features into $y$’s coordinate frame, thereby realizing directional, embedding-level message passing throughout the network.
Cellular sheaves and copresheaves: contrasting graph learning
While much of the recent literature has focused on sheaf learning (Ayzenberg et al., 2025), our approach is rooted in the copresheaf perspective. Copresheaves offer a novel perspective on deep learning by assigning each vertex $x$ a feature space $\mathcal{F}(x)$, and each edge $x\to y$ a linear map $\rho_{x\to y}:\mathcal{F}(x)\to\mathcal{F}(y)$ that governs data propagation. This setup departs from the traditional deep learning’s core assumption of a single, shared latent space, enabling the modeling of heterogeneous, task-specific latent spaces and directional relations. The significance of this approach lies in its ability to generalize and connect different deep learning paradigms. Copresheaf-type architectures extend beyond sheaf neural networks (Hansen and Gebhart, 2020b; Bodnar et al., 2022b; Barbero et al., 2022c; Duta et al., 2024; Battiloro et al., 2024a) and topological neural network architectures (Papillon et al., 2023b)—typically designed for non-Euclidean data—by also accommodating Euclidean data effectively. This versatility allows them to unify applications across diverse data domains and architectural frameworks, providing a cohesive structure that leverages directional information flow and adapts to task-specific requirements.
@@
The data on nodes $x$ and $y$, denoted by $\mathbf{h}_{x}\in\mathcal{F}(x)$ and $\mathbf{h}_{y}\in\mathcal{F}(y)$, “agree” on the edge $e$ if their images under the restriction maps coincide:
@@
A global section of a sheaf on a graph $G$ assigns data $\mathbf{h}_{v}\in\mathcal{F}(v)$ to each vertex $v$ and $\mathbf{h}_{e}\in\mathcal{F}(e)$ to each edge $e$, such that for every edge $e$ between nodes $x$ and $y$: $\mathcal{F}_{x\triangleleft e}(\mathbf{h}_{x})=\mathcal{F}_{y\triangleleft e}(\mathbf{h}_{y})$. This consistency condition ensures data consistency across local connections in a network. Global sections represent equilibrium states where this local agreement holds across the entire graph, enabling unified and coherent data representations for the consensus of a complex system. Most sheaf-based architectures have focused on diffusion-type models (Hansen and Gebhart, 2020b; Bodnar et al., 2022b; Barbero et al., 2022c; Duta et al., 2024; Battiloro et al., 2024a), where the sheaf Laplacian $\Delta_{\mathcal{F}}$ minimizes the Dirichlet energy, thereby enforcing global consistency condition. More precisely, for some arbitrary choice of orientation for each edge, define the coboundary map
$\delta:C^{0}\to C^{1},\qquad(\delta\mathbf{h})_{e}=\mathcal{F}_{x\triangleleft e}(\mathbf{h}_{x})\;-\;\mathcal{F}_{y\triangleleft e}(\mathbf{h}_{y})\quad\text{for }e=(x,y),$ (2)
@@
Unlike sheaves, which enforce data consistency across overlaps, copresheaves model directional data flow—ideal for processes like information propagation, causality, and hierarchical dependencies. They assign vector spaces $F(x)$ only to vertices and define learnable linear maps $\rho_{x\to y}\colon\mathcal{F}(x)\to\mathcal{F}(y)$ along directed edges, without imposing sheaf consistency conditions. This vertex-centric, anisotropic framework naturally integrates with message-passing methods such as GNNs and TNNs, allowing parameterized maps to adapt during training.

### v3

@@ Abstract
We introduce copresheaf topological neural networks (CTNNs), a powerful unifying framework that encapsulates a wide spectrum of deep learning architectures, designed to operate on structured data, including images, point clouds, graphs, meshes, and topological manifolds. While deep learning has profoundly impacted domains ranging from digital assistants to autonomous systems, the principled design of neural architectures tailored to specific tasks and data types remains one of the field’s most persistent open challenges. CTNNs address this gap by formulating model design in the language of copresheaves, a concept from algebraic topology that generalizes most practical deep learning models in use today. This abstract yet constructive formulation yields a rich design space from which theoretically sound and practically effective solutions can be derived to tackle core challenges in representation learning, such as long-range dependencies, oversmoothing, heterophily, and non-Euclidean domains. Our empirical results on structured data benchmarks demonstrate that CTNNs consistently outperform conventional baselines, particularly in tasks requiring hierarchical or localized sensitivity. These results establish CTNNs as a principled multi-scale foundation for the next generation of deep learning architectures.
@@ 1 Introduction
Deep learning has excelled by exploiting structural biases, such as convolutions for images (Krizhevsky et al., 2012), transformers for sequences (Vaswani et al., 2017), and message passing for graphs (Gilmer et al., 2017). However, the design of architectures that generalize across domains with complex, irregular, or multiscale structure remains a notorious challenge (Bronstein et al., 2017; Hajij et al., 2023b). Real-world data, which span physical systems, biomedical signals, and scientific simulations, rarely adhere to the regularity assumptions embedded in conventional architectures. These data are inherently heterogeneous, directional, and hierarchical, often involving relations beyond pairwise connections or symmetric neighborhoods.
@@
Convolutional neural networks (CNNs), designed for uniform grids, do not fully capture local irregularities; graph neural networks (GNNs) often rely on homophily and tend to oversmooth feature representations as depth increases; and transformers, while excellent at capturing long-range dependencies, assume homogeneous embedding spaces, incur quadratic complexity, and lack built-in notions of anisotropy or variable local structures. These shortcomings highlight the need for a framework that can natively encode diverse local behaviors, respect directional couplings, and propagate information across scales without sacrificing local variations or imposing unwarranted homogeneity.
@@
To address this foundational gap, we propose copresheaf topological neural networks (CTNNs), a unifying framework for deep learning based on copresheaves, a categorical structure that equips each local region of a domain with its own feature space, along with learnable maps specifying how information flows between regions. Unlike traditional models that assume a global latent space and isotropic propagation, our framework respects local variability in representation and directional flow of information, enabling architectures that are multiscale, anisotropic and expressive.
@@
Figure 1: A copresheaf topological neural network (CTNN) operates on combinatorial complexes ($C$s), which generalize Euclidean grids, graphs, meshes, and hypergraphs. A CTNN is characterized by a set of locally indexed copresheaf maps $\rho_{x_{i}\to x_{j}}$, defined between cells $x_{i}$ and $x_{j}$ in the CC, and directed from $x_{i}$ to $x_{j}$. The figure illustrates how a CTNN updates a local representation $\mathbf{h}_{x}$ of a cell $x$ using neighborhood representations $\mathbf{h}_{y}$ and $\mathbf{h}_{z}$, which are sent to $x$ via the learnable local copresheaf maps $\rho_{x\to x}$, $\rho_{y\to x}$ and $\rho_{z\to x}$.
@@
By constructing CTNNs on combinatorial complexes (CCs) (Hajij et al., 2023b, a), which generalize graphs, simplicial complexes, and cell complexes, we enable a principled message-passing mechanism over general topological domains formulated within the theory of copresheaves. This unified perspective subsumes many deep learning paradigms, including GNNs, attention mechanisms, sheaf neural networks (Hansen and Ghrist, 2019b; Bodnar et al., 2022), and topological neural networks (Papillon et al., 2023; Hajij et al., 2023b; Bodnar et al., 2021a; Ebli et al., 2020; Giusti et al., 2023) within a single formalism. Our approach further departs from the traditional assumption of a single shared latent space by modeling task-specific, directional latent spaces that bridge diverse deep learning frameworks. Furthermore, CTNNs flexibly handle both Euclidean and non-Euclidean data, supporting expressive architectures such as copresheaf GNNs, transformers, and CNNs, which learn structure-aware, directional transport maps. CTNNs offer a promising framework for developing next-generation, topologically informed, structure-aware machine learning models. See Figure 1 for an illustration.
2 Related Work
@@
Our work here is related to sheaf neural networks (SNNs), which extend traditional GNNs by employing the mathematical framework of cellular sheaves to capture higher-order or heterogeneous relationships. Early work by Hansen and Ghrist (2019a, b) introduced methods to learn sheaf Laplacians from smooth signals and developed a spectral theory that connects sheaf topology with graph structure. Building on these ideas, Hansen and Gebhart (2020) proposed the first SNN architecture, demonstrating that incorporating edge-specific linear maps can improve performance on tasks involving asymmetric or heterogeneous relations.
@@
Recent advances have focused on mitigating common issues in GNNs, such as oversmoothing and heterophily. For instance, Bodnar et al. (2022) introduced neural sheaf diffusion processes that address these challenges by embedding topological constraints into the learning process. Similarly, Barbero et al. (2022a, b) developed connection Laplacian methods and attention-based mechanisms that further enhance the expressiveness and efficiency of SNNs. The versatility of the sheaf framework has also been demonstrated through its extension to hypergraphs and heterogeneous graphs (Duta et al., 2023; Braithwaite et al., 2024), which enables modeling of higher-order interactions. Moreover, novel approaches incorporating joint diffusion processes (Hernandez Caralt et al., 2024) and Bayesian formulations (Gillespie et al., 2024) have improved the robustness and uncertainty quantification of SNNs. Finally, the application of SNNs in recommender systems (Purificato et al., 2023) exemplifies their practical utility in real-world domains. Together, these contributions demonstrate the potential of SNNs to enrich graph-based learning by integrating topological and geometric information directly into neural architectures. Our proposed CTNNs generalize these architectures while avoiding restrictive co-boundary maps or rank-specific Laplacian operators. Appendix I provides a more thorough literature review of related work.
@@ 3 Preliminaries
This section presents preliminary concepts needed for developing our theoretical framework. It revisits CCs and neighborhood structures, reviews sheaves and copresheaves on directed graphs, and compares cellular sheaves with copresheaves in graph-based modeling.
@@ 3.1 Combinatorial Complexes and Neighborhood Structures
To ensure generality, we base our framework on CCs (Hajij et al., 2023b, a), which unify set-type and hierarchical relations over which data are defined. CC-neighborhood functions then formalize local interactions forming a foundation for defining sheaves and higher-order message passing schemes for CTNNs.
@@
A CC is a triple $(\mathcal{S},\mathcal{X},\operatorname{rk})$, where $\mathcal{S}$ is a finite non-empty set of vertices, $\mathcal{X}\subset\mathcal{P}(\mathcal{S})\setminus\{\emptyset\}$, with $\mathcal{P}(\mathcal{S})$ denoting the power set of $\mathcal{S}$, and $\operatorname{rk}:\mathcal{X}\to\mathbb{Z}_{\geq 0}$ is a rank function such that if $\{s\}\in\mathcal{X}$, $\operatorname{rk}(\{s\})=0$ for all $s\in\mathcal{S}$, and $x\subseteq y\implies\operatorname{rk}(x)\leq\operatorname{rk}(y)$ for all $x,y\in\mathcal{X}$.
@@
When context permits, we write a CC $(\mathcal{S},\mathcal{X},\operatorname{rk})$ simply as $\mathcal{X}$. Each $x\in\mathcal{X}$ has rank $\operatorname{rk}(x)$, and $\dim\mathcal{X}=\max_{x\in\mathcal{X}}\operatorname{rk}(x)$. We refer to elements of $\mathcal{X}$ by cells. The $k$-cells $x^{k}$ of $\mathcal{X}$ are defined to be the cells $x$ with $\operatorname{rk}(x)=k$. We use the notation $\mathcal{X}^{k}=\{x\in\mathcal{X}:\operatorname{rk}(x)=k\}=\operatorname{rk}^{-1}(\{k\})$. See Fig 2 for an example.
@@
A neighborhood function on a CC $(\mathcal{S},\mathcal{X},\operatorname{rk})$ is a map $\mathcal{N}:\mathcal{X}\to\mathcal{P}(\mathcal{X})$, which assigns to each cell $x$ in $\mathcal{X}$ a collection of neighbor cells $\mathcal{N}(x)\subset\mathcal{X}$, referred to as the neighborhood of $x$. In our context, two neighborhood functions are commonly used, namely the adjacency neighborhood $\mathcal{N}_{\mathrm{adj}}(x)=\{\,y\in\mathcal{X}\mid\exists\,z\in\mathcal{X}:x\subset z,\;y\subset z\}$ and the incidence neighborhood $\mathcal{N}_{\mathrm{inc}}(x)=\{\,y\in\mathcal{X}\mid x\subset y\}$.
@@
Let $\mathcal{N}$ be a neighborhood function on a CC $\mathcal{X}$. Let $\mathcal{Y}=\{y_{1},\dots,y_{n}\}\subset\mathcal{\mathcal{X}}$, $\mathcal{Z}=\{z_{1},\dots,z_{m}\}\subset\mathcal{\mathcal{X}}$ be two collections of cells such that $\mathcal{N}(y_{j})\subseteq\mathcal{Z}$ for all $1\leq j\leq n$. An element of the neighborhood matrix $\mathbf{G}\in\{0,1\}^{m\times n}$ is defined as
[…]
@@ 3.2 Sheaves and Copresheaves on Directed Graphs
The copresheaf formalism assigns each vertex its unique feature space $\mathcal{F}(x)$, respecting the potentially heterogeneous nature of the data, and each directed edge a transformation $\rho_{x\!\to\!y}$ that tells how data move between those spaces, $\mathcal{F}(x)\to\mathcal{F}(y)$. This separation between where data reside and how they travel provides a foundation for learning beyond the single-latent-space assumption of standard deep learning architectures. Concretely, copresheaves are defined as follows [^1] While we avoid overly complicated jargon, the appendix links our constructs to those of category theory for a more rigorous exposition.:
Definition 4 (Copresheaf on directed graphs).
@@
Think of a copresheaf as a system for sending messages across a network, where each node has its own language (stalk), and edges translate messages (linear maps) to match the recipient’s language. More specifically, on a directed graph $G=(V,E)$, each vertex $x\in V$ carries a task-specific latent space $\mathcal{F}(x)$, and every edge $x\!\to\!y\in E$ applies a learnable, edge-indexed linear map $\rho_{x\to y}:\mathcal{F}(x)\to\mathcal{F}(y)$ that re-embeds $x$’s features into $y$’s coordinate frame, thus realizing directional, embedding-level message passing throughout the network.
While much of the recent literature has focused on sheaf learning (Ayzenberg et al., 2025), our approach is based on a copresheaf perspective. This setup departs from the traditional deep learning core assumption of a single, shared latent space, enabling the modeling of heterogeneous, task-specific latent spaces and directional relations. The significance of this approach lies in its ability to generalize and connect different deep learning paradigms. Copresheaf-type architectures extend beyond SNNs (Hansen and Gebhart, 2020; Bodnar et al., 2022; Barbero et al., 2022b; Duta et al., 2023; Battiloro et al., 2024b) and TNN architectures (Hajij et al., 2023b), typically designed for non-Euclidean data, by also accommodating Euclidean data effectively. This versatility allows them to unify applications across diverse data domains and architectural frameworks, providing a unified structure that uses directional information flow and adapts to task-specific requirements.
@@
For any cell $c$ (vertex or edge), the vector space $\mathcal{F}(c)$ is typically called the stalk at $c$. The data on nodes $x$ and $y$, denoted by $\mathbf{h}_{x}\in\mathcal{F}(x)$ and $\mathbf{h}_{y}\in\mathcal{F}(y)$, “agree” on the edge $e$ if their images under the restriction maps coincide:
@@
A global section of a sheaf on a graph $G$ assigns data $\mathbf{h}_{v}\in\mathcal{F}(v)$ to each vertex $v$ and $\mathbf{h}_{e}\in\mathcal{F}(e)$ to each edge $e$, such that for every edge $e$ between nodes $x$ and $y$ it holds that $\mathcal{F}_{x\triangleleft e}(\mathbf{h}_{x})=\mathcal{F}_{y\triangleleft e}(\mathbf{h}_{y})$. This consistency condition ensures data consistency across local connections in a network. Global sections represent equilibrium states, where this local agreement holds across the entire graph, enabling unified data representations for a complex system. Most sheaf‐based architectures have focused on diffusion‐type models (Hansen and Gebhart, 2020; Bodnar et al., 2022; Barbero et al., 2022b; Duta et al., 2023; Battiloro et al., 2024b), where the sheaf Laplacian $\Delta_{\mathcal{F}}$ minimizes the Dirichlet energy, ensuring global consistency. Precisely, let $C^{0}(G;\mathcal{F})=\bigoplus_{v\in V}\mathcal{F}(v)$ and $C^{1}(G;\mathcal{F})=\bigoplus_{e\in E}\mathcal{F}(e)$ denote the spaces of vertex-valued and edge-valued cochains of the sheaf $\mathcal{F}$, respectively. Then for some arbitrary choice of orientation for each edge, define the coboundary map
$\delta:C^{0}(G;\mathcal{F})\to C^{1}(G;\mathcal{F}),\qquad(\delta\mathbf{h})_{e}=\mathcal{F}_{x\triangleleft e}(\mathbf{h}_{x})-\mathcal{F}_{y\triangleleft e}(\mathbf{h}_{y})\quad\text{for }e=(x,y)\in E,$ (2)
@@
Unlike sheaves, which ensure data consistency across overlaps, copresheaves model directional data flow, making them well-suited for processes such as information propagation, causality, and hierarchical dependencies. Copresheaves assign vector spaces $\mathcal{F}(x)$ only to vertices and define learnable linear maps $\rho_{x\to y}:\mathcal{F}(x)\to\mathcal{F}(y)$ along directed edges, without imposing sheaf consistency constraints. This vertex-centric, anisotropic framework naturally integrates with message-passing architectures such as GNNs and TNNs, allowing parameterized maps to adapt during training. See Appendix B.3 for further discussion, and Appendix C for the definition and properties of the copresheaf Laplacian.